OpenStudy (jazzyluvsu123):

Super entrepreneur Freddie Fandango thinks his latest product might set the world on fire. It's basically a giant microwave that cooks lots of whole hams and turkeys in just a few seconds. If the microwave was 9 feet long, 6 feet wide, and 3 feet high, what is its volume in cubic yards?

OpenStudy (lochana):

\[volume = height.width.long\]

OpenStudy (lochana):

and 1 yard = 3feets

OpenStudy (lochana):

do you want me to solve?

OpenStudy (jazzyluvsu123):

So 9*3, 6*3, 3*3 then 27*18*9= my answer?

OpenStudy (lochana):

no. don't multiply by 3. divide it

OpenStudy (jazzyluvsu123):

hmm okai 9/3= 3 6/3=2 3/3=1 Then What?

OpenStudy (lochana):

9 feets in yards = 9/3 = 3yards 6 feets in yards = 9/6 = 1.5yards 3 feets in yards = 9/3 = 1yards volume = 3*1.5*1yards^2 = 4.5 cubic yards

OpenStudy (jazzyluvsu123):

o.o

OpenStudy (lochana):

i am sorry,

OpenStudy (lochana):

9 feets in yards = 9/3 = 3yards 6 feets in yards = 6/3 = 2yards 3 feets in yards = 3/3 = 1yards volume = 3*2*1yards^3 = 6 cubic yards

OpenStudy (jazzyluvsu123):

Got it ! Thanks

OpenStudy (lochana):

you are welcome:)